OpenStudy (anonymous):

The train from Metropolis A to Metropolis B takes 4 hours and 15 minutes. If the rail distance between Metropolis A and Metropolis B is 250 miles, how fast must the train average? 58.8 55.8

OpenStudy (anonymous):

58.8

OpenStudy (mathmale):

This is a "rate" problem: (How many miles)/(How many hours = how many miles per hour.) \[\frac{ 250 miles }{4hr 15 \min }=\frac{ 250 miles }{ 4.25 hours }=58.4 mph.\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (mathmale):

Excuse me, that's 58.8 mph. Units of measurement are important here.